OpenStudy (anonymous):

BaCl2 (aq) + K2SO4 (aq) → BaSO4 (s) + 2 KCl (aq) What would you multiply "moles of barium chloride" by to convert to the units "moles of potassium chloride" ?

OpenStudy (anonymous):

Explain how ?

OpenStudy (frostbite):

We can set up a relationship between the species in your reaction equation: For a reaction a A + b B -> c C + d D it then follow: \[\large \frac{ n(A) }{ a }=\frac{ n(B) }{ b }=\frac{ n(C) }{ c }=\frac{ n(D) }{ d }\] Try use the relation and tell me how we could convert then.

OpenStudy (anonymous):

How can I used that ?